A student wanted to find the mass of calcium carbonate in an indigestion tablet. She crushed up a tablet and added an excess of HCL acid (25cm3 of 1.00 mol/dm3). She the titrated the excess against 0.5 mol/dm3 of NaOH requiring 25.8cm3 of the NaOH. Calculate the mass off calcium carbonate in the tablet.
1
Expert's answer
2017-11-01T14:03:54-0400
Ca〖CO〗_3+2HCl→CaCl_2+H_2 O+〖CO〗_2
Excess of HCl: HCl+NaOH→NaCl+H_2 O
ν(HCl)=c∙V=0.025∙1=0.025 mol.
ν(NaOH)=c∙V=0.0258∙0.5=0.0129 mol=〖ν(HCl)〗_excess.
〖ν(HCl)〗_(for Ca〖CO〗_3 )=0.025-0.0129=0.0121 mol.
ν(Ca〖CO〗_3 )=0.0121/2=0.00605 mol.
M(Ca〖CO〗_3 )=100.1 g/mol.
m(Ca〖CO〗_3 )=0.00605∙100.1=0.605605 g≈606 mg.